# Diffusion MRI with Machine Learning

### v1

@@ III-D Data, ground truth, and evaluation
Most commonly, however, high-quality dMRI measurements have been used to generate the data needed for developing and validating machine learning methods. For tract segmentation/registration, Li et al. used an automatic method based on probabilistic tractography and atlas-based segmentation to generate their training labels [253]. For micro-structure mapping and for FOD estimation, published studies have either used dense q-space data and estimated ground truth parameters using standard methods (e.g., [73, 119, 120, 126, 118, 126]) or they have used simulation techniques such as the Monte Carlo method (e.g., [43, 85, 67]). Despite its limitations, data simulation is sometimes the only reasonable approach when high-quality data are not available to compute a reliable ground truth. For micro-structure estimation, both numerical and physical phantom are considered valid and powerful methods for developing and validating computational methods [127, 372]. Data simulation also has potential advantages: (i) it enables exploring the full range of parameters that may influence the signal, and (ii) the ground truth target is known accurately [67, 96]. Because of the flexibility of data simulation methods, they may also be useful for investigating what data acquisition protocols lead to more accurate reconstruction. This approach was used by Lee et al. to determine the optimal b values for IVIM reconstruction with deep learning [373]. Gyori et al. investigated the impact of training data distribution on the test performance of supervised machine learning methods for micro-structure mapping [67]. When training on simulated data, they found that in order for the model to achieve reasonable accuracy on atypical brain tissue, the training data should be synthesized using the complete range of plausible parameter values. When the parameter space used to generate the training data matched that of healthy brains, estimation accuracy on atypical brains was low, although the estimations showed a deceptively high precision.
@@
There have been innovative approaches to dMRI data simulation in prior works. Nedjati et al. used Monte Carlo simulation to synthesize dMRI data based on histologically valid micro-structure parameters [25]. Ye and Li have developed a method for synthesizing q-space data that can potentially be used for any machine learning application in dMRI [374]. Their signal generator is an MLP that is trained to minimize the difference between the distribution of the synthesized signal and that of the observed signal. They use a continuous representation of the dMRI data in the SHORE basis [375]. Their preliminary experiments have shown promising results for estimation of the NODDI model. Qin et al. have proposed and validated a knowledge transfer method that relies on a source domain where high-quality data are available [376]. Using the SHORE basis, this method interpolates the source data in the q-space and/or voxel space to synthesize data that match the quality of the target data. Model training is carried out in the source domain and transferred to the target domain. For Gibbs artifact suppression, one study trained a CNN on more than one million natural images and simulated more than 10 million data instances [332]. Experiments showed that the size and richness of the collection of natural images used for training was essential to ensuring generalizability of the model to dMRI test data. Graham et al. developed a simulation method to enable direct quantitative evaluation of artifact-correction methods [377], which they subsequently used to develop and validate a deep learning-based motion artifact detection technique [334]. Their work showed that with proper data simulation, only a small real dataset was need to calibrate the method.
@@
Using simulated data is typically much more convenient than real-world data. Some works have relied solely on simulated data to train and or evaluate their methods [88, 32, 15, 53]. However, using in-silico simulations to synthesize training data may be inadequate because some of the factors such as the noise distribution are difficult to model accurately [85]. For IVIM-DKI estimation, one study observed that increasing the noise level in the synthesized training data improved the model’s test accuracy [87]. Masutani [96] analyzed the impact of matching the noise level between training and test data. He found that the best test performance was achieved when the noise level in the training data was close to or slightly higher than the noise level in the test data. Martins et al. [36], on the other hand, found that training with a much noisier dataset could still result in superior results on test data with relatively lower noise. For DKI estimation, another work found that varying the noise strength in the synthetic training data led to better results on clinical test scans compared with a model that was trained using a single noise strength [97]. In some applications, such as cardiac DTI [32] and fetal imaging [74] that we have mentioned above, obtaining reliable in-vivo data faces especial difficulties such that using synthetic training data may be unavoidable.
@@
Another important consideration is the choice of evaluation metrics. It is often unclear what metrics should be reported to present a comprehensive assessment of the method and a fair comparison among competing techniques. As an example, for micro-structure mapping most studies report the RMSE as the main or the only metric. However, it has been suggested that correlation analysis and sensitivity analysis are important for revealing the limitations of machine learning-based methods that cannot be assessed based solely on estimation error [68]. Moreover, rather than reporting a single global RMSE value for a method, detailed analysis of the estimation error such as its variations in different brain regions may reveal biases of machine learning methods that are not captured by global RMSE [378, 66, 67, 53].
@@
Although for many dMRI computations obtaining the ground truth may be difficult or impossible, test-retest reproducibility assessments may be far easier to perform. As a matter of fact, repeatability tests have often been used to assess machine learning methods in dMRI. For example, Tunc et al. used repeated scans of the same subjects to assess the repeatability of their tract segmentation method [211], while Barbieri et al. computed inter-subject variability to validate a deep learning method for IVIM estimation [35]. Performing such tests has become more feasible with the growing availability of public datasets that include repeated scans of the same subjects on the same scanner [379] or on different scanners [380] or on different scanners with different protocols [381, 293].
@@
Where ground truth is difficult to establish, visual assessment by medical experts may be a viable alternative. One study quantified the agreement between two readers in their interpretation of IVIM parameter maps to compare a deep learning method with conventional estimation methods [35]. However, the authors have pointed out that this approach may favor methods that wrongly compute consistent results while failing to account for genuine heterogeneity. Another work has reported successful validation of a deep learning method for fetal DTI estimation based on blind randomized assessment and scoring by neuroanatomists and neuroradiologists [74]. Experts may also be able to help in generating ground truth labels for training and validation. This has been a common approach for white matter tract segmentation. This approach, however, may suffer from difficulty of defining anatomical ground truth [37], high intra- and inter-expert variability [156], and labeling errors. Such label errors need to be properly accounted for using advanced machine learning techniques [268]. In applications such as tractography and tract analysis, obtaining labels from multiple experts may be useful but costly [173, 22, 151].
@@
A useful approach to assessing the potential of machine learning for dMRI analysis is via open competitions, where research teams are free to develop and apply standard non-learning methods and machine learning techniques. A good example is a recent open challenge where research teams were invited to develop deep learning methods for estimating diffusion tensor parameters from reduced number of measurements [77]. Specifically, focusing on distinguishing between episodic and chronic migraine patients, the goal was to investigate whether deep learning methods were able to achieve the same level of sensitivity and specificity when using 21 measurements, compared with 61 measurements typically used with standard methods for this analysis. With conventional estimation methods, 60% of the differences detected by TBSS with 61 measurements are missed when using 21 measurements. A total of 14 research teams participated in the challenge. Results showed that deep learning methods improved the global image metrics such as PSNR and SSIM. They also improved sensitivity of TBSS. However, these improvements came at the cost of higher false positive rates, which increased linearly with the true positive rate. The teams had used different deep learning models and had adopted various approaches, such as mapping the reduced measurements directly to the target DTI parameters or estimating dense dMRI measurements first. The study concluded that the results of deep learning methods should be interpreted with caution even when global image quality metrics are improved.
As we mentioned above, an appealing approach to assessing the new methods is to use measures of patient outcome as the evaluation target. This approach has been successfully attempted in a few works, where the predicted micro-structure indices have been correlated with post‐stroke outcome [70] and outcome of pancreatic cancer patients [34]. Another study used acute-phase dMRI (as well as anatomical MRI) data as input to a small neural network to directly predict the chronic size of ischemic lesions, where the T2 images at 3 months after stroke were used to establish the ground truth [382]. This method achieved an Area Under the Receiver Operator Characteristic curve (AUROC) of 0.89.
III-E Explainability, uncertainty, and reliability
@@
The significance of these issues for machine learning-based dMRI analysis is clear. Even though very few studies have investigated these issues, they have reported important results. For example, it has been reported that machine learning methods display deceptively confident (i.e., high-precision) predictions on noisy test data, while with classical estimation methods higher measurement noise levels are properly reflected in predictions as higher variance [67]. Unfortunately, very little work has been devoted to characterize these issues and to devise effective solutions. A few studies have attempted to incorporate uncertainty estimation in their methods. For micro-structure mapping, Ye et al. quantified the uncertainty of their neural network-based micro-structure estimation using a residual bootstrap technique [64]. They validated their uncertainty computation method by assessing the correlation with estimation error and inspecting the confidence intervals, and they concluded that this method could compute practically useful uncertainty estimates. For diffusion tensor estimation, another work computed data-dependent uncertainty via loss attenuation and model uncertainty using Monte Carlo dropout [378]. Their extensive experiments showed that estimation uncertainties computed by their proposed methods could highlight the model’s biases, detect domain shift, and reflect the strength of noise in the measurements. Avci et al. have shown that a simple uncertainty quantification method based on dropout can reduce the prediction error and also serve as an indicator of pathologies or measurement artifacts [393]. For IVIM estimation, one study proposed to estimate the full Gaussian posteriors (i.e., mean and variance) for the parameters and to use the standard deviation of the posterior as a proxy for estimation uncertainty [89]. Experiments with simulated and real data showed that the computed uncertainties were qualitatively meaningful. For tract segmentation, one work has used dropout and test-time augmentation to compute epistemic and aleatoric uncertainties [194]. After applying post-hoc uncertainty calibration, the computed uncertainty is shown to be well correlated with segmentation error. The authors argue that such uncertainty computations can be useful in surgical planning applications such as for reliable delineation of eloquent areas. In general, it may be easier to probe the explainability of conventional machine learning models than large neural networks. For example, for micro-structure estimation, one study was able to extensively examine the importance of different feature sets and different measurement shells in the q-space for a random forest model [21]. Performing a similar analysis for a deep learning model should be more challenging due to the longer training times for these models.
@@
In the context of dMRI super-resolution, Tanno et al. estimated the uncertainty for CNN and random forest models [282, 394]. For a CNN estimator, they employed a heteroscedastic noise model to compute the intrinsic uncertainty and variational dropout to compute the model uncertainty [395, 282]. However, their evaluation of estimation uncertainty relied on a qualitative visual inspection of the correlation between uncertainty and estimation error. They showed that the uncertainty was higher for brain pathologies not observed in the training data and that the computed estimation uncertainty could be useful for downstream computations such as tractography. In another work [394], they employed similar techniques to compute the estimation uncertainty for dMRI super-resolution based on the IQT approach discussed above. They also developed methods to propagate the uncertainty estimates to downstream computations such as computation of micro-structure indices. Through experiments they showed that uncertainty quantification could highlight model’s failures, explain its prediction performance, and improve prediction accuracy. Also in the context of super-resolution, Qin et al. [283] used a method based on model ensembles [396] to quantify the uncertainty for estimation of NODDI parameters. They showed that the computed uncertainties had moderate correlation with estimation error and could be used to improve estimation accuracy.
@@
Model interpretability has also received little attention. Xu et al. [181] used attention maps similar to [397] in order to interpret the decision mechanism of a tractography-based tract segmentation model. They showed that the attention maps provided useful insights and were consistent with the anatomy of the fiber tracts connecting different brain regions. Varadarajan and Haldar criticized nonlinear machine learning FOD estimation methods for their black-box nature and difficulty of predicting their performance on unseen measurement schemes [398]. Instead, they formulated FOD computation as a linear estimation problem and achieved prediction accuracy on par with or better than standard methods. One can argue that works of Ye et al. that involve unfolding the iterative algorithms for micro-structure estimation [63, 61] also represent a small but significant step towards improving the explainability of deep learning models. Rensonnet et al. [344] have further improved Ye’s work in this direction by proposing a hybrid method that involves the use of a physics-based fingerprinting dictionary within a deep learning framework. Their method first projects the dMRI signal into a dictionary containing physics-based responses. An MLP network then uses the sparse representations computed in the first stage to compute the target micro-structure indices. Compared with an end-to-end trained MLP, this new method showed competitive accuracy while having qualitatively more interpretable features.
@@
Better model explainability may also aid in designing more effective computational methods. For example, it is unclear why machine learning methods often achieve more accurate micro-structure estimation than standard optimization-based methods. The shortcomings of optimization methods are due to several factors including sensitivity to measurement noise, poor initialization, and degeneracy (i.e., presence of multiple equally-optimal solutions). For DTI estimation, which consists in a linear or non-linear least squares problem, the advantage of deep learning models are likely only due to their ability to use the measurements in neighboring voxels to reduce the impact of measurement noise. For multi-compartment models with a complex optimization landscape, on the other hand, machine learning models may actually be able to learn to avoid local minima and compute better estimates.
IV Conclusions
@@
Machine learning has the potential to usher in a new generation of computational techniques for dMRI processing and analysis. Some of the capabilities that have been demonstrated by multiple studies include: estimation of micro-structural indices, tractography and tract analysis tasks such as segmentation and registration, denoising, artifact correction, and data harmonization and enhancement. Compared with conventional computational techniques in dMRI, the new machine learning methods may offer improved performance in several aspects including: faster computation, ability to handle imperfect data, more flexible modular design as well as end-to-end optimization, seamless and easy integration of spatial information and other sources of information such as anatomical MRI. However, in order to realize the full potential of these methods, we need to overcome several critical limitations and remaining issues. More rigorous validation on rich and heterogeneous datasets in necessary. Standardized data pre-processing pipelines, validation benchmarks, and evaluation metrics can dramatically help the research community in developing and assessing more effective methods and discovering truly meritorious techniques. Finally, enabling model explainability and proper uncertainty estimation may greatly enhance and expedite the adoption of these methods in clinical and scientific applications.
V Acknowledgements
@@
This research was supported in part by the National Institute of Neurological Disorders and Stroke and Eunice Kennedy Shriver National Institute of Child Health and Human Development of the National Institutes of Health (NIH) under award numbers and R01HD110772 R01NS128281. The content of this publication is solely the responsibility of the author and does not necessarily represent the official views of the NIH.

### v2

@@ V-D4 Data simulation
Another approach is to use simulation techniques such as the Monte Carlo method. This approach has been widely used for tissue microstructure mapping [60, 178, 168]. Despite its limitations, data simulation is sometimes the only reasonable approach when high-quality data are not available to compute a reliable ground truth. For microstructure estimation, numerical simulations and physical phantoms are considered valid and powerful methods for developing and validating computational methods [211, 396]. Data simulation has two potential advantages: (i) it enables exploring the full range of parameters that may influence the signal, and (ii) the ground truth target is known accurately [168, 167]. Because of the flexibility of data simulation methods, they may also be useful for investigating what data acquisition protocols lead to more accurate reconstruction. This approach was used by Lee et al. to determine the optimal b values for IVIM reconstruction with deep learning [397]. Gyori et al. investigated the impact of training data distribution on the test performance of supervised machine learning methods for microstructure mapping [168]. When training on simulated data, they found that in order for the model to achieve reasonable accuracy on atypical brain tissue, the training data should be synthesized using the complete range of plausible parameter values. When the parameter space used to generate the training data matched that of healthy brains, estimation accuracy on atypical brains was low, although the estimations showed a deceptively high precision.
@@
There have been innovative approaches to dMRI data simulation in prior works. Nedjati et al. used Monte Carlo simulation to synthesize dMRI data based on histologically valid microstructure parameters [45]. Ye and Li have developed a method for synthesizing q-space data that can potentially be used for any machine learning application in dMRI [398]. Their signal generator is an MLP that is trained to minimize the difference between the distribution of the synthesized signal and that of the observed signal. They use a continuous representation of the dMRI data in the SHORE basis [122]. Their preliminary experiments have shown promising results for estimation of the NODDI model.
Qin et al. have proposed and validated a knowledge transfer method that relies on a source domain where high-quality data are available [399]. Using the SHORE basis, this method interpolates the source data in the q-space and/or voxel space to synthesize data that match the quality of the target data. Model training is carried out in the source domain and transferred to the target domain. For Gibbs artifact suppression, one study trained a CNN on more than one million natural images and simulated more than 10 million data instances [90]. Experiments showed that the size and richness of the collection of natural images used for training was essential to ensuring generalizability of the model to dMRI test data. Graham et al. developed a simulation method to enable direct quantitative evaluation of artifact-correction methods [400]. They subsequently used this method to develop and validate a deep learning motion artifact detection technique [92]. Their work showed that with proper data simulation only a small real dataset was sufficient to calibrate the method.
@@
Using simulated data is typically much more convenient than real-world data. Some works have relied solely on simulated data to train and or evaluate their methods [181, 50, 33, 147]. However, using in-silico simulations to synthesize training data may be inadequate because some factors such as the noise distribution are difficult to model accurately [178]. For IVIM-DKI estimation, one study observed that increasing the noise level in the synthesized training data improved the model’s test accuracy [180]. Masutani [167] analyzed the impact of matching the noise level between training and test data. He found that the best test performance was achieved when the noise level in the training data was close to or slightly higher than the noise level in the test data. Martins et al. [21], on the other hand, found that training with a much noisier dataset could still result in superior results on test data with relatively lower noise. For DKI estimation, another work found that varying the noise level in the synthetic training data led to better results on clinical test scans compared with a model that was trained using a single noise level [401]. In some applications, such as cardiac DTI [50] and fetal imaging [22] mentioned above, obtaining reliable in-vivo data faces especial difficulties such that using synthetic training data may be unavoidable.
@@
Another important consideration is the choice of evaluation metrics. It is often unclear what metrics should be reported for a comprehensive assessment of a new method and a fair comparison with competing techniques. As an example, for tissue microstructure mapping most studies report the RMSE as the main or the only metric. However, it has been suggested that correlation analysis and sensitivity analysis are important for revealing certain limitations of machine learning methods that cannot be assessed based solely on RMSE [169]. Moreover, rather than reporting a single global RMSE value for a method, detailed analysis of the estimation error such as its variation in different brain regions may reveal biases that are not captured with global RMSE [402, 164, 168, 147].
@@
Although for many dMRI computations obtaining the ground truth may be difficult or impossible, test-retest reproducibility assessments may be far easier to perform. In fact, repeatability tests have often been used to assess machine learning methods in dMRI. For example, Tunc et al. used repeated scans of the same subjects to assess the repeatability of their tract segmentation method [293], while Barbieri et al. computed inter-subject variability to validate a deep learning method for IVIM estimation [53]. Performing such tests has become more feasible with the growing availability of public datasets that include repeated scans of the same subjects on the same scanner [403] or on different scanners [404] or on different scanners with different protocols [40, 41].
@@
Where an objective/quantitative ground truth is difficult to establish, visual assessment by human experts may be a viable alternative. One study quantified the agreement between two readers in their interpretation of IVIM parameter maps to compare a deep learning method with conventional estimation methods [53]. However, the authors have pointed out that this approach may favor methods that wrongly compute consistent results while failing to account for genuine heterogeneity. Another work has reported successful validation of a deep learning method for fetal DTI estimation based on blind randomized assessment and scoring by neuroanatomists and neuroradiologists [22].
@@
An effective approach to assessing the potential of machine learning for dMRI analysis is via open competitions, where research teams are free to develop and apply standard non-learning methods and machine learning techniques. A good example is a recent open challenge where research teams were invited to develop deep learning methods for estimating diffusion tensor parameters from reduced measurements [173]. Specifically, focusing on distinguishing between episodic and chronic migraine patients, the goal was to investigate whether deep learning methods were able to achieve the same level of sensitivity and specificity when using 21 measurements, compared with 61 measurements typically used with standard methods for this analysis. With conventional estimation methods, 60% of the differences detected by TBSS with 61 measurements are missed when using 21 measurements. A total of 14 research teams participated in the challenge. Results showed that deep learning methods improved the global image metrics such as PSNR and SSIM. They also improved the sensitivity of TBSS. However, these improvements came at the cost of higher false positive rates, which increased linearly with the true positive rate. The teams had used different deep learning models and had adopted various approaches, such as mapping the reduced measurements directly to the target DTI parameters or estimating dense dMRI measurements first. The study concluded that the results of deep learning methods should be interpreted with caution even when global image quality metrics are improved.
A very appealing approach to assessing the new methods is to use measures of patient outcome as the evaluation target. This approach has been successfully attempted in a few works, where the predicted microstructure indices have been correlated with post‐stroke outcome [161] and outcome of pancreatic cancer patients [52]. Another study used acute-phase dMRI (as well as anatomical MRI) data as input to a small neural network to directly predict the chronic size of ischemic lesions, where the T2-weighted images at 3 months after stroke were used to establish the ground truth [405].
V-E Explainability, uncertainty, and reliability
@@
Even though very few studies have investigated these issues for machine learning-based dMRI analysis, they have reported important observations. For example, it has been reported that machine learning methods produce deceptively confident (i.e., high-precision) predictions on noisy test data, while with classical estimation methods higher measurement noise levels are properly reflected in predictions as higher variance [168]. Unfortunately, very little work has been devoted to characterize these issues and to devise effective solutions.
@@
A few studies have incorporated uncertainty estimation in their methods. For microstructure mapping, Ye et al. quantified the uncertainty of their neural network-based tissue microstructure estimation using a residual bootstrap technique [160]. They validated their uncertainty computation method by assessing the correlation with estimation error and inspecting the confidence intervals. They concluded that this method could compute practically useful uncertainty estimates. For diffusion tensor estimation, another work computed data-dependent uncertainty via loss attenuation and model uncertainty via Monte Carlo dropout [402]. Their extensive experiments showed that estimation uncertainties computed by their proposed methods could highlight the model’s biases, detect domain shift, and reflect the measurement noise level. Avci et al. have shown that a simple uncertainty quantification method based on dropout reduces the prediction error and can also serve as an indicator of pathology or measurement artifacts [416]. For IVIM estimation, one study estimated the full Gaussian posteriors (i.e., mean and variance) for the parameters and used the standard deviation of the posterior as a proxy for estimation uncertainty [182]. Experiments with simulated and real data showed that the computed uncertainties were qualitatively meaningful. For tract segmentation, one work has used dropout and test-time augmentation to compute epistemic and aleatoric uncertainties [305]. After applying post-hoc uncertainty calibration, the computed uncertainty is shown to be well correlated with segmentation error. The authors argue that such uncertainty computations can be useful in surgical planning applications such as for reliable delineation of eloquent areas. In general, it may be easier to probe the explainability of conventional machine learning models than large neural networks. For example, for microstructure estimation, one study was able to extensively examine the importance of different feature sets and different measurement shells in the q-space for a random forest model [42]. Performing a similar analysis for a deep learning model should be more challenging due to the longer training times for these models.
@@
In the context of dMRI super-resolution, Tanno et al. estimated the uncertainty for CNN and random forest models [111, 417]. For their CNN estimator, they employed a heteroscedastic noise model to compute the intrinsic uncertainty and variational dropout to compute the model uncertainty [111]. However, their evaluation of estimation uncertainty relied on a qualitative visual inspection of the correlation between uncertainty and estimation error. They showed that the uncertainty was higher for brain pathologies not observed in the training data and that the computed estimation uncertainty could be useful for downstream computations such as tractography. In another work [417], they employed similar techniques to compute the estimation uncertainty for dMRI super-resolution based on the IQT approach discussed above. They also developed methods to propagate the uncertainty estimates to downstream analyses such as computation of microstructure indices. Experiments showed that uncertainty quantification could highlight model’s failures, explain prediction performance, and improve prediction accuracy. Also in the context of super-resolution, Qin et al. [105] used a method based on model ensembles [418] to quantify the uncertainty for estimation of NODDI parameters. They showed that the computed uncertainties had moderate correlation with estimation error and could be used to improve estimation accuracy.
@@
Model interpretability has also received little attention. Xu et al. [270] used attention maps (similar to [419]) to interpret the decision mechanism of a tractography-based tract segmentation model. They showed that the attention maps provided useful insights that were consistent with the anatomy of the fiber tracts connecting different brain regions. For subject-level classification based on tract-specific tissue microstructure, one study has shown that class activation maps can identify the white matter tracts that underlie the group differences [420]. Varadarajan and Haldar criticized nonlinear machine learning FOD estimation methods for their black-box nature and difficulty of predicting their performance on unseen measurement schemes [421]. They formulated FOD computation as a linear estimation problem and achieved prediction accuracy on par with or better than standard methods. One can argue that works of Ye et al. that involve unfolding the iterative algorithms for microstructure estimation [159, 158] also represent a step towards improving the explainability of deep learning models. Rensonnet et al. [368] have further improved Ye’s work in this direction by proposing a hybrid method that involves the use of a physics-based fingerprinting dictionary within a deep learning framework. Their method first projects the dMRI signal into a dictionary containing physics-based responses. Subsequently, an MLP uses the sparse representations computed in the first step to compute the target microstructure indices. Compared with an MLP that is trained end-to-end, this new method showed competitive accuracy while having qualitatively more interpretable features.
@@
Better model explainability may also aid in designing more effective computational methods. For example, it is unclear why machine learning methods often achieve more accurate tissue microstructure estimation than standard optimization-based methods. The shortcomings of optimization methods are due to several factors including sensitivity to measurement noise, poor initialization, and degeneracy (i.e., presence of multiple equally-optimal solutions). For DTI estimation, which consists in a linear or non-linear least squares problem, the advantage of deep learning models are likely only due to their ability to use the measurements in neighboring voxels to reduce the impact of measurement noise. For multi-compartment models with a complex optimization landscape, on the other hand, machine learning models may actually be able to learn to avoid local minima and compute better estimates.
VI Conclusions
@@
Machine learning has the potential to usher in a new generation of computational techniques for dMRI data processing and analysis. Some of the capabilities that have been demonstrated by multiple studies include: denoising, artifact correction, data harmonization, estimation of microstructural indices, tractography, and tract analysis tasks such as segmentation and registration. Compared with conventional computational techniques in dMRI, the new machine learning methods may offer improved performance in several aspects including: faster computation, ability to handle imperfect data, flexible modular design, end-to-end optimization, seamless and easy integration of spatial information and other sources of information such as anatomical MRI. However, in order to realize the full potential of these methods, we need to overcome several critical limitations and remaining issues. More rigorous validation on rich and heterogeneous datasets in necessary. Standardized data preprocessing pipelines, validation benchmarks, and evaluation metrics can dramatically help the research community in developing and assessing more effective methods and discovering truly meritorious techniques. Finally, enabling model explainability and proper uncertainty estimation may facilitate and expedite the adoption of these methods in clinical and scientific applications.
Funding
@@
This research was supported in part by the National Institute of Neurological Disorders and Stroke and Eunice Kennedy Shriver National Institute of Child Health and Human Development of the National Institutes of Health (NIH) under award numbers R01HD110772 and R01NS128281. The content of this publication is solely the responsibility of the author and does not necessarily represent the official views of the NIH.
Declaration of Competing Interests
The author has no conflicts of interest to declare.